### v2

@@ Abstract
Coxeter pointed out that a number of polytopes can be projected orthogonally into two dimensions in such a way that their vertices lie on a number of concentric regular triacontagons (or 30-gons). Among them are the 600-cell and 120-cell in four dimensions and Gosset’s polytope $4_{21}$ in eight dimensions. We show how these projections can be modified into Kochen-Specker diagrams from which parity proofs of the Bell-Kochen-Specker theorem are easily extracted. Our construction trivially yields parity proofs of fifteen bases for all three polytopes and also allows many other proofs of the same type to be constructed for two of them. The defining feature of these proofs is that they have a fifteen-fold symmetry about the center of the Kochen-Specker diagram and thus involve both rays and bases that are multiples of fifteen. Any proof of this type can be written as a word made up of an odd number of distinct letters, each representing an orbit of fifteen bases. Knowing a word makes it possible to write down all the characteristics of the associated proof without first having to recover its bases. A comparison is made with earlier approaches that have been used to obtain parity proofs in these polytopes, and two questions related to possible applications of these polytopes are raised.
@@ 1 Introduction
Coxeter pointed out that a number of four and eight dimensional polytopes can be projected orthogonally into two dimensions in such a way that their vertices lie on a number of concentric regular triacontagons (or 30-gons). He dubbed this the triacontagonal projection and the frontispiece of his monograph Regular PolytopesCox1 shows such a projection of the 600-cell. His book Introduction to GeometryCox2 shows a similar projection of the 120-cell and his paper with Shephardleonardo shows the projection of $4_{21}$ [^1] Interestingly, hand drawn versions of all three projections existed before the computer versions were made. According to Coxeter, the projection of the 600-cell was drawn by van Oss, that of the 120-cell by Chilton and, even earlier, by Wythoff and that of $4_{21}$ by Peter McMullen.. All three projections have $C_{30}$ point symmetry, which gives them a highly symmetrical appearance. Table 1 lists the features of these polytopes that are of interest in connection with proofs of the Bell-Kochen-Specker (BKS) theoremBKS KS  [^2] This theorem rules out the existence of deterministic noncontextual hidden variables theories as an alternative to quantum mechanics., namely, the system of rays and bases they give rise to. A ray is an undirected line through the center of the polytope that passes through a pair of antipodal vertices, either of which may be taken as a representative of the ray. The number of rays is half the number of vertices of the polytope. A basis is a set of four or eight mutually orthogonal rays (four for the 600-cell and 120-cell and eight for Gossets’s polytope).
@@
Table 1: The names and symbols of the three polytopes studied in this paper are listed in the first column, their dimension in the second and their ray-basis symbol in the third. The ray-basis symbol lists the number of rays to the left of the dash and the number of bases to the right, with the subscript on the left indicating the number of bases to which each ray belongs and that on the right the number of rays in each basis.
@@
The purpose of this paper is to show how the triacontagonal projections of these polytopes can be modified into Kochen-Specker diagrams from which parity proofs of the BKS theorem are easily extracted. It is interesting that these projections, which arose out of purely geometrical considerations, should lead so directly to an important physical result.
The plan of this paper is as follows. The next section spells out the simple procedure that turns a triacontagonal projection into a Kochen-Specker diagram and points out the forms that the rays and bases of the polytope take in it, particularly their fifteen-fold symmetry about its center. The notion of a parity proof is introduced and its relationship to the basis table is explained. We do this to make it easier for readers with no knowledge of the BKS theorem to follow the chain of arguments leading from Coxeter’s projections to parity proofs, the objects of interest in this paper. The three sections that follow discuss how parity proofs can be extracted for each of the polytopes. Since the construction is basically the same in all the cases, we describe it in detail only for the 120-cell and discuss only the individual peculiarities and results obtained in the other two cases. The final section compares the approach used here with others that have been used to find parity proofs in these polytopes and comments on its advantages and limitations relative to them. We end by raising two questions suggested by this work, to which we do not have answers.
@@
Before proceeding, we say a few words about the field of contextuality studies in which this work is situated. The basic result that laid the foundations for the field was the discovery by BellBKS and Kochen and SpeckerKS of the incompatibility between deterministic noncontextual hidden variables theories and quantum mechanics. Bell arrived at this conclusion by a continuum argument, whereas Kochen and Specker established it using a finite set of rays in a three-dimensional Hilbert space. The latter approach has been greatly generalized over the years, and many examples of rays are now known in real and complex Hilbert spaces of all dimensions that prove the theorem (see Budroni for many references to such work). The three polytopes studied in this paper are just one of the many settings in which such proofs arise. What makes them particularly interesting is that they each have well over a billion parity proofs in them, in addition to a vast number of other contextuality proofs that are not as easy to characterize.
The notion of noncontextuality has evolved considerably over the years as a result of questions about how effectively it can be tested through experiments of limited precision Exp -Barrett . In a notable development, Spekkens et. al.Spekkens1 have provided an operationally meaningful definition of a noncontextual theory in terms of the predictions it makes for the most general types of preparations and measurements that can be carried out on a quantum system. The advantage of this approach is that it makes it possible to design experiments in which noncontextual theories predict results only within a restricted range of parameters, whereas quantum mechanics (and/or experiment) yield results outside that range. Two examples of processes for which this can be done are parity-oblivious multiplexingSpekkens2 and universal quantum computation via magic state distillationHoward . In both cases, the “quantum advantage” obtained by implementing the process with quantum resources can be related to the breakdown of noncontextuality (and hence the relevance of contextuality) as an explicative mechanism for it.
Kochen-Specker sets have also been shown to be useful for a variety of other tasks such as ”all versus nothing” proofsLiu , bipartite perfect quantum strategiesBPQS , the quantum computational advantage of shallow circuitsBrayvi and random number generationrandom . For all these reasons, the study of quantum contextuality remains an active area of research.
@@
2 From the triacontagonal projection to the Kochen-Specker diagram
@@
The triacontagonal projections show the vertices of the polytopes as points and their edges as line segments joining the points. Two changes must be made to convert them into Kochen-Specker diagrams [^3] The Kochen-Specker diagram of a system of rays is a graph whose vertices are the rays and whose edges join orthogonal pairs of rays.: firstly, every diametrically opposite pair of points in the projection must be replaced by just one of them and, secondly, the line segments representing the edges of the polytope must be replaced by line segments joining orthogonal pairs of its rays. The first change is needed because diametrically opposite pairs of points in the projection correspond to antipodal pairs of vertices of the polytope, which define the same ray. This change can be effected by keeping only the alternate vertices of all the triacontagons, thereby converting them into pentadecagons (or 15-gons). The second change can be effected by joining all pairs of pentadecagon vertices that correspond to orthogonal rays (this is easily done by using the triacontagonal coordinates of the polytopes [^4] The triacontagonal coordinates of the 600-cell and 120-cell are given in Cox1 and those of Gosset’s polytope can be found in Waegell2015 .).
@@
The resulting Kochen-Specker diagram differs in a number of ways from the triacontagonal projection: it has only half the number of points, a very different pattern of lines and a fifteen-fold symmetry about its center (instead of the thirty-fold symmetry of the original). However its visual appearance is confusing and not particularly helpful, so we will replace it by a table of numbers, the so called “basis table”, that captures all the information in it. The basis table is just a list of all the bases formed by the rays of the polytope. Since all the polytopes we are considering are saturated (i.e., they have no orthogonalities between their rays other than those represented in their bases), the basis table captures all the orthogonalities in the Kochen-Specker diagram and serves as a convenient substitute for it.
The Kochen-Specker diagrams we will be dealing with, which are all derived from triacontagonal projections, have the special feature that both their rays and their bases have a fifteen-fold symmetry about the center of the diagram. The symmetry of the rays is evident, but that of the bases less so. However it follows from the fact that fifteen-fold rotations about the center of the Kochen-Specker diagram that send it into itself correspond to four-dimensional rotations of the polytope that send it into itself [^5] This is most easily seen from the triacontagonal coordinates of the polytopes.. We will exploit this symmetry by introducing a numbering scheme for the rays that greatly simplifies the specification of the bases. This turns out to be crucial because the bases in two of the polytopes number in the hundreds or thousands and a direct specification of them would be impractical.
Our numbering scheme is as follows. The rays of any polytope are numbered in counterclockwise order around its pentadecagons, starting from the outermost pentadecagon and proceeding inwards, with the lowest number on any pentadecagon being attached to the ray/vertex lying on, or just above, the positive x-axis. The fifteen-fold symmetry of the bases about the center of the pentadecagons then makes it necessary to specify only one basis in each orbit, because the others can be obtained by repeatedly adding one to each of its rays, provided that one goes back to the ray at the beginning of a pentadecagon when one overshoots the one at the end (a procedure we will refer to as “wraparound”). We term the representative basis we pick from each orbit the “generator” and label it by a lower case letter. The basis table can then be expressed as a “word” made up of a number of distinct letters, each representing a generator. The entire basis table can be easily reconstructed from the word representing it.
A parity proof of the BKS theorem [^6] To get a better feeling for what a parity proof is, go to the website https://quantumgames.wpi.edu/ and click on the tab “600-cell” to see many examples of such proofs in this polytope. can be obtained from a basis table by picking an odd number of bases in such a way that each of the rays that occurs in them occurs in an even number of them. [^7] The reason this proves the theorem is that it is impossible to assign a 0 or a 1 to the rays in a noncontextual fashion in such a way that each basis has exactly one ray assigned the value 1 in it.. The parity proofs of interest in this paper are those having a fifteen-fold symmetry about the center of the Kochen-Specker diagram and, as such, can be represented by “words” made up of an odd number of distinct letters representing generators (and their orbits of fifteen bases). How such proofs can be picked out in the three polytopes is a task that we will address in the next three sections.
@@
We end with a word about terminology. Since the parity proofs of this paper have a fifteen-fold symmetry, they might be termed “pentadecagonal” proofs. However we have referred to them as “triacontagonal” proofs in the title of this paper to acknowledge their origin in the projections of the same name due to Coxeter.
@@ 3 The 600-cell
The 600-cell is a four-dimensional regular polytope with 120 vertices lying on the surface of a 3-sphere. Its vertices come in antipodal pairs, so it has 60 rays. The rays form 75 bases, with each ray occurring in five bases, so the ray-basis symbol of this polytope is $60_{5}-75_{4}$.
The triacontagonal projection of the 600-cell consists of 120 points representing its vertices and 720 line segments representing its edges. The Kochen-Specker diagram obtained from it has 60 points representing the rays and 450 line segments connecting orthogonal pairs of rays. The rays lie, in sets of fifteen, at the vertices of four concentric regular pentadecagons; see Table 2 for the numbering of the rays and their polar coordinates in the plane of the projection [^8] Expressions for the circumradii of the pentadecagons can be given in terms of radicals but they are not simple, particularly for the two polytopes studied later. Thus we content ourselves by giving numerical values for the circumradii in Tables 2,5 and 10. The expressions for the circumradii can be found in the references mentioned in footnote 4..
Pentadecagon Circumradius Angle Rays
A 1.0000 0 1-15
B 0.8135 6 16-30
C 0.6728 6 31-45
D 0.3383 0 46-60
Table 2: Pentadecagons of the 600-cell, labeled A through D in order of decreasing circumradius. The rays, which lie at the vertices of the pentadecagons, are numbered in counter-clockwise order around them, beginning from the outermost pentadecagon and proceeding inwards. The ray with the lowest number on each pentadecagon is the one lying on or just above the positive x-axis. The rays on A and D begin from the x-axis while those on B and C begin from $6{{}^{\circ}}$ above the x-axis. The rays on any pentadecagon are spaced $24{{}^{\circ}}$ apart.
The 75 bases of the 600-cell fall into five orbits of fifteen bases each. Table 3 gives one generator for each orbit, with its basis profile shown in the column to its left; the four letters of the profile indicate the pentadecagons to which the rays of the generator (as well as all the bases in its orbit) belong. The first two generators have unique profiles, whereas the last three have a common profile. However, the last three generators have orbits that are distinct and have no bases in common. The two polytopes we study later also have this feature, i.e., they have generators with the same profile that nevertheless have no bases in common.
@@
Table 3: Generators of the 600-cell. The generators are shown in the second column, with each being labeled by a lower case letter. The first column shows the basis profiles of the generators, with each capital letter indicating the pentadecagon to which the corresponding ray of the generator belongs. The first two generators have unique profiles, whereas the last three all have a common profile. The orbits of these generators are shown in Table 4.

### v3

@@ Abstract
Coxeter pointed out that a number of polytopes can be projected orthogonally into two dimensions in such a way that their vertices lie on a number of concentric regular triacontagons (or 30-gons). Among them are the 600-cell and 120-cell in four dimensions and Gosset’s polytope $4_{21}$ in eight dimensions. We show how these projections can be modified into Kochen-Specker diagrams from which parity proofs of the Bell-Kochen-Specker theorem are easily extracted. Our construction trivially yields parity proofs of fifteen bases for all three polytopes and also allows many other proofs of the same type to be constructed for two of them. The defining feature of these proofs is that they have a fifteen-fold symmetry about the center of the Kochen-Specker diagram and thus involve both rays and bases that are multiples of fifteen. Any proof of this type can be written as a word made up of an odd number of distinct letters, each representing an orbit of fifteen bases. Knowing the word representing a proof makes it possible to infer all its characteristics without first having to recover its bases. A comparison is made with earlier approaches that have been used to obtain parity proofs in these polytopes, and some directions in which this work can be extended are discussed.
@@ 1 Introduction
Coxeter pointed out that a number of four and eight dimensional polytopes can be projected orthogonally into two dimensions in such a way that their vertices lie on a number of concentric regular triacontagons (or 30-gons). He dubbed this the triacontagonal projection and the frontispiece of his monograph Regular PolytopesCox1 shows such a projection of the 600-cell. Similar projections of the 120-cell and Gosset’s polytope $4_{21}$ are shown in Cox2 and leonardo , respectively. All three projections have $C_{30}$ point symmetry about their center, which gives them a highly symmetrical appearance. Table 1 lists the features of these polytopes that are of interest in connection with proofs of the Bell-Kochen-Specker (BKS) theoremBKS ; KS  [^1] This theorem rules out the existence of deterministic noncontextual hidden variables theories as an alternative to quantum mechanics. See Sec.2 for an overview of it., namely, the system of rays and bases they give rise to. A ray is an undirected line through the center of the polytope that passes through a pair of antipodal vertices, either of which may be taken as a representative of the ray. The number of rays of each of the polytopes, whose vertices come in antipodal pairs, is half the number of its vertices. A basis is a set of four or eight mutually orthogonal rays (four for the 600-cell and 120-cell and eight for Gossets’s polytope).
@@
Table 1: The three polytopes studied in this paper, with their Schläfli symbols (for the first two) or Coxeter symbol (for the third). The dimension of the polytope is indicated in the second column and its ray-basis symbol in the third. The number to the left of the dash in the symbol is the number of rays of the polytope and the number to the right is the number of bases, with the subscript on the left indicating the number of bases to which each ray belongs and that on the right the number of rays in each basis.
@@
The purpose of this paper is to show how the triacontagonal projections of the three polytopes can be converted into Kochen-Specker diagrams from which parity proofs of the BKS theorem are easily extracted. It is interesting that these projections, which arose out of purely geometrical considerations, should lead so directly to an important physical result.
The plan of this paper is as follows. Sec.2 gives a quick introduction to the BKS theorem and the notion of a parity proof that should allow readers unfamiliar with these topics to follow the rest of the (largely geometrical) arguments in this paper. Sec.3 introduces the triacontagonal projection and shows how it can be modified into a Kochen-Specker diagram from which parity proofs of the BKS theorem can be extracted. A Kochen-Specker diagram (or orthogonality graph) is a graph whose vertices are rays and whose edges connect orthogonal pairs of rays. An important feature of a Kochen-Specker diagram obtained from a triacontagonal projection is that both the rays and the bases [^2] This will become clear in Sec.3. have a fifteen-fold symmetry about the center of the diagram. As a result, it has parity proofs with a fifteen-fold symmetry about the center of the diagram and it is only with such proofs that this paper will be concerned. We will refer to these proofs as “fifteen-fold proofs” or simply proofs in the rest of this paper. There are a large number of other proofs of lower symmetry, or no symmetry at all, but we will largely ignore them except as they relate to the proofs discussed here.
All parity proofs in the three polytopes considered here have been found earlier by a variety of other methods Waegell2011b ; Waegell2014 ; Waegell2015 ; Pavicic2011 ; Lisonek2014b , with our fifteen-fold proofs included among them. However, the symmetry of the proofs went unrecognized and was not exploited in any way. We show that by recognizing the symmetry at the outset, it is possible to categorize the proofs better and also give a more detailed account of the members of particular subclasses. In particular, we show that the proofs of highest symmetry in all three polytopes can be written as “words” made up of a number of distinct letters from which all their properties can be inferred.
Sections 4,5 and 6 discuss the specific form the triacontagonal projection takes for each of the polytopes and how it facilitates the extraction of the fifteen-fold parity proofs in them. The task turns out to be trivial for the 600-cell, which has only a handful of such proofs, but is more involved in the other two cases in which the number of such proofs runs into the billions. We use an approach due to Lisoněk et allisonek1 , suitably modified to take advantage of the symmetry of the triacontagonal projection, to show how the proofs can be written as “words” made up of distinct letters, each of which represents an orbit of fifteen bases. We give a simple algorithm for constructing all the words representing arbitrary fifteen-fold parity proofs of the 120-cell and Gosset’s polytope and show how all the characteristics of the proofs can be inferred from the words themselves without the need to recover the bases involved in the proof. Since many of the proofs consist of hundreds of bases, the benefits of such an approach are obvious.
Section 7 recapitulates the main results of the paper and also indicates some directions in which this work might be extended.
@@
Before going on, we give a brief overview of the field of quantum contextuality in which the present work is situated. The basic result that laid the foundations for the field was the discovery by BellBKS and Kochen and SpeckerKS of the incompatibility between deterministic noncontextual hidden variables theories and quantum mechanics. Bell arrived at this conclusion by a continuum argument, but Kochen and Specker established it by using 117 rays in a three-dimensional Hilbert space. The latter approach has been greatly generalized over the years, and many examples are now known of rays in both real and complex Hilbert spaces of all dimensions that prove the theorem (see Budroni for many references to such work). The three polytopes studied in this paper are just one of the many venues in which such proofs have been found. What makes them particularly interesting is that they each have well over a billion parity proofs in them, in addition to a vast number of other contextuality proofs that are not as easy to categorize or recognize.
Quantum contextuality has been shown to be useful for a variety of tasks such as “all versus nothing” proofsLiu , bipartite perfect quantum strategiesBPQS , the quantum computational advantage of shallow circuitsBrayvi , random number generationrandom and universal quantum computation via magic state distillationHoward . In a notable development, Spekkens Spekkens1 broadened the notion of a noncontextual theory by specifying the sorts of restrictions that would govern its predictions about the most general sorts of measurements that could be carried out on arbitrarily prepared systems. The advantage of such an approach is that it makes it possible to identify tasks in which a noncontextual theory predicts outcomes only within a certain range of parameters whereas experiment (and contextual models) allow values outside that range. This approach makes it possible to demonstrate the existence of quantum contextuality by means of robust experiments that are not subject to the sort of objections that had been made earlier to traditional contextuality testsMeyer Clifton . An example of a practical task in which a “quantum advantage” can be demonstrated by using Spekkens’ approach is parity oblivious multiplexingSpekkens2 .
2 The Bell-Kochen-Specker (BKS) theorem
We give a quick introduction to the BKS theorem and the notion of a parity proof for readers unfamiliar with these topics so that they can follow the rest of the paper (see Budroni for a more detailed treatment of this topic and many references).
The BKS theorem rules out the existence of deterministic noncontextual hidden variables theories as an alternative to quantum mechanics. To understand what such theories are, consider a set of rays in a Hilbert space of even dimension $N$ that form a number of bases, with each basis consisting of $N$ mutually orthogonal rays from the set. It is possible for a ray to occur in more than one basis of the set (and in the ray-basis sets of interest to us, which are shown in the third column of Table 1, this always happens). A noncontextual hidden variables theory is one that assigns the value 0 or 1 to every ray in the set, and further requires every ray to assume its assigned value in every basis in which it occurs. A contextual theory, on the other hand, allows a ray to switch its value between 0 and 1 in different bases.
It turns out that it is possible to do a laboratory experiment to determine whether each ray in a basis has the value 0 or 1, and that such an experiment always shows exactly one ray in any basis to have the value 1, with all the others having the value 0. In order to agree with this empirical finding, a noncontextual theory must be able to assign a 0 or 1 to every ray in a ray-basis set in such a way that every basis has just one ray assigned the value 1 in it. However, if there is a ray-basis set in an even number of dimensions with the properties that (a) the number of bases in the set is odd, and (b) each ray occurs in an even number of bases of the set, then the noncontextual theory cannot succeed in its task and must be rejected. The argument showing why it fails is this: every basis is required to have exactly a single 1 in it, and from (a) the total number of 1’s over all the bases must be odd. However, from (b), the total number of 1’s over all the bases must be even, and this contradicts the earlier conclusion. Because this argument relies on an even/odd contradiction, the ray-basis set it uses to achieve its goal is often referred to as a parity proof.
The purpose of the BKS theorem is to rule out the existence of noncontextual hidden variables theories, and one way of doing this is by finding a ray-basis set that satisfies the conditions stated above for a parity proof. That is not the only way of proving the theorem, but it is the method of choice for the polytopes considered in this paper.
@@ 2 The Bell-Kochen-Specker (BKS) theorem
3 From the triacontagonal projection to the Kochen-Specker diagram and the basis table
The triacontagonal projection is an orthogonal projection of certain polytopes from their original dimension down to two dimensions in such a way that their vertices lie on a number of concentric regular triacontagons. Coxeter constructed such projections for the three polytopes studied here by obtaining triacontagonal coordinates for their vertices [^3] The triacontagonal coordinates of the 600-cell and 120-cell are given in Cox1 and those of Gosset’s polytope in leonardo . with the property that the first two coordinates specify the positions of the vertices in the plane of the projection and the remaining coordinates their positions in the subspace orthogonal to the projection plane.
@@ 3 From the triacontagonal projection to the Kochen-Specker diagram and the basis table
The triacontagonal projection shows the vertices of the polytope as points and its edges as line segments joining the points. Two changes must be made to convert it into a Kochen-Specker diagram (or orthogonality graph): firstly, every diametrically opposite pair of points in the projection should be replaced by just one of them and, secondly, the line segments representing edges should be replaced by line segments joining orthogonal pairs of rays. The first change is needed because diametrically opposite triacontagon vertices in the projection correspond to antipodal vertices of the polytope, both of which represent the same ray. This change can be effected by keeping only the alternate vertices of all the triacontagons, thereby converting them into pentadecagons (or 15-gons). The second change can be effected by using the triacontagonal coordinates of the vertices to work out all the orthogonalities between the rays and then looking at just the first two coordinates to pick out all orthogonal pairs of rays in the projection plane.
@@
The Kochen-Specker diagram differs from its parent projection in having only half the number of points, a very different pattern of lines and a fifteen-fold symmetry about its center compared to the thirty-fold symmetry of the parent. However the visual appearance of the diagram is confusing and not particularly helpful, and we will make no direct use of it. Instead, we will replace it by a table of numbers, the so called basis table, that captures all the essential information in it. The basis table is simply a list of all the bases formed by the rays of the polytope. Since all the polytopes we are considering are saturated (i.e., have no orthogonalities between rays other than those contained in their bases), the basis table captures all the orthogonalities in the Kochen-Specker diagram and so serves as a complete substitute for it. However the basis table is more than just a convenient proxy for the Kochen-Specker diagram. As will become clear below, it facilitates the search for parity proofs, and also their viewing, in ways that the diagram simply cannot.
We will exploit the fifteen-fold symmetry of the Kochen-Specker diagram by introducing a numbering scheme for the rays that greatly simplifies the specification of the bases. We will number the rays in counterclockwise order around the pentadecagons, beginning from the outermost one and proceeding inwards towards smaller ones, always attaching the smallest number on any pentadecagon to the ray lying on, or just above, the positive x-axis. Then a rotation by one-fifteenth of a revolution (or $24^{\circ}$) about the center of the diagram sends any ray into the one whose number is one greater than it, with ray 1 going into ray 2, ray 24 into ray 25, and so on. However, a modification of this rule is clearly needed for rays that are multiples of fifteen, like 30 or 75, because such rays do not increase by one but go back into the rays at the beginning of their pentadecagons instead; thus ray 30 goes back into ray 16, ray 75 back into ray 51, and so on. We will term this drop in the number of a ray as one goes around a pentadecagon as “wraparound” (see Table 4, and its caption, for a graphic illustration of this concept as one runs one’s eye down the columns).
Next pick a set of rays in the diagram that form a basis [^4] This is easily done using the triacontagonal coordinates of the rays.. If one adds one to the numbers of each of the rays in this basis (with wraparound!), one can verify, using triacontagonal coordinates, that the new rays so obtained also form a basis. This pattern holds no matter how often it is repeated and shows that the bases have a fifteen-fold symmetry about the center of the diagram, just like the rays. In view of this, it is possible to generate an orbit of fifteen bases by picking any one of its members and repeatedly adding one to each of its rays (with wrapround). We will term the representative picked from each orbit the generator and label it by a lower case letter. The entire basis table can then be written as a “word” made up of a number of distinct letters, each representing a generator, and all the bases can be recovered by replacing each letter (or generator) by the orbit of fifteen bases it represents.
A fifteen-fold parity proof must consist of a number of complete orbits of bases (to respect its fifteen-fold symmetry), and so can be written as a word made up of the letters/generators representing its orbits of bases. Since the total number of bases in a parity proof must be odd, and each letter contributes fifteen bases, the word representing a fifteen-fold proof must consist of an odd number of letters. However not all odd letter words are suitable because each ray is required to occur an even number of times over the bases. To see which odd letter words are suitable, it is convenient to introduce a capital letter for each ray in a basis to indicate the pentadecagon to which it belongs. Then every lower case letter in a word can be replaced by the set of capital letters representing the pentadecagons that cover its bases and the word can be written as a string of capital letters, each representing a complete pentadecagon’s worth of fifteen rays. But the requirement that each ray occur an even number of times over the word implies that each capital letter also occur an even number of times. Thus the condition that an odd letter word must satisfy to be a parity proof can be stated as follows: replace each lower case letter in the word by its set of capital letters and count the number of times each capital letter occurs over the word; if all the counts are even the word represents a parity proof, otherwise not. A concrete illustration of this procedure is given in Eq.(1) for the 600-cell and Eq.(3) for the 120-cell, with the present explanation repeated there and supplemented by the details needed in the individual cases to make it clear.
In the next three sections we will describe an efficient method of picking out odd length words representing parity proofs in all three polytopes.
@@
We end with a word about terminology. Since the parity proofs of this paper have a fifteen-fold symmetry, they might best be termed “pentadecagonal” proofs. However we have referred to them as “triacontagonal” proofs in the title of this paper to acknowledge their origin in the projections of the same name due to Coxeter.
@@ 4 The 600-cell
The 600-cell is a four-dimensional regular polytope with 120 vertices lying on the surface of a 3-sphere. Its vertices come in antipodal pairs, so it has 60 rays. The rays form 75 bases [^5] While the bases can be calculated using any of the standard coordinates for the vertices, the advantage of using triacontagonal coordinates is that it allows the points corresponding to the bases to be picked out in the projection plane by looking at just the first two coordinates of the rays., with each ray occurring in five bases, so the ray-basis symbol of this polytope is $60_{5}-75_{4}$.
The triacontagonal projection of the 600-cell has 120 points representing its vertices and 720 line segments representing its edges. The Kochen-Specker diagram obtained from it has 60 points representing the rays and 450 line segments connecting orthogonal pairs of rays. The rays lie, in sets of fifteen, at the vertices of four concentric regular pentadecagons. Table 2 lists the numbering scheme we use for the rays.
Pentadecagon Rays
A 1-15
B 16-30
C 31-45
D 46-60
Table 2: Pentadecagons of the 600-cell, labeled A through D in order of decreasing circumradius. The rays, which lie at the vertices of the pentadecagons, are numbered in counter-clockwise order around them, with the lowest number on any pentadecagon being attached to the ray lying on or just above the positive x-axis, and the numbers increasing as one moves inwards towards smaller pentadecagons. The pentadecagons A and D have vertices lying on the x-axis, whereas B and C have vertices lying $6{{}^{\circ}}$ above the x-axis.
The 600-cell has 75 bases that fall into five orbits of fifteen bases each. Table 3 gives a generator for each orbit along with its basis profile, with each capital letter of the profile indicating the pentadecagon to which the corresponding ray of the generator belongs.
@@
Table 3: Generators of the 600-cell. The five generators are labeled $a$ through $e$ and the four mutually orthogonal rays of which each is made up are indicated after it in curly brackets. The basis profile of each generator is indicated in the column to its left, with each capital letter indicating the pentadecagon to which the corresponding ray of the generator belongs. The first two generators have unique profiles, whereas the last three have a common profile. Despite this, the generators $c$,$d$ and $e$ give rise to orbits that have no bases in common (as is clear from Table 4).